A chemist wants to mix a 60% saline solution with 20 liters of a 45% saline solution to obtain a 55% saline solution. How much of the 60% solution should she use?

Answer by checkley77(12844)

.80*20+.45x=.55(20+x)
16+.45x=11+.55x
.45x-.55x=11-16
-.10x=-5
x=-5/-.1
x=50 liters of 60% solution is used.
Proof:
.80*20+.45*50=.55(20+50)
16+22.5=.55*70
38.5=38.5